If I have a center of (1,1), a focus of (1,3) and a vertex of (1,5) , what is a^2 and b^2 in my equation equal to?
Answer by lwsshak3(11628)

If I have a center of (1,1), a focus of (1,3) and a vertex of (1,5) , what is a^2 and b^2 in my equation equal to?
**
This is an ellipse with vertical major axis.
Its standard form of equation: (x-h)^2/b^2+(y- k)^2/a^2=1, (h,k)=(x,y) coordinates of center
..
center: (1,1) (given)
a=4 (from 1 to 5)
a^2=16
..
c=2 (from 1 to 3)
c^2=4
..
c^2=a^2-b^2
b^2=a^2-c^2=16-4=12
..
ans:
a^2=16
b^2=12
..
Equation: (x-1)^2/12+(y-1)^2/16=1